MA113 Information for instructors Spring 2001 Course coordinator: Russell Brown Textbook: The textbook for this course will be Calculus, 3rd edition, by James Stewart, ISBN 0-534-21798-2 Syllabus: A course calendar with suggested problems is available. You may duplicate the version I have provided to Elizabeth. Each instructor should also provide an individual syllabus giving name, contact information and office hours and additional information about grading policy. The course calendar and a copy of my syllabus are available at http://www.ms.uky.edu/∼rbrown/ma113cc as tex documents. I expect that each instructor will prepare an additional page that details grading policies and gives their name, office hours etc. The University Ombudsperson has made several requests for syllabi in all courses. Please see the ombud’s memo for the full story.

1. The grading procedure for the course should be described. Instructors who plan to curve grades are asked to describe their curving procedure in the syllabus.

2. Instructors who plan to take attendance should consider defining what is meant by absence in the syllabus.

Math resource center: The mathematics resource center is scheduled to open this fall in the basement of the White Hall Classroom Building. Until the center is open, we will operate temporarily out of the Inslab, CB313. All teaching assistants are asked to schedule at least one of their office hours in this facility (and a total of three office hours). I would like to staff the block of time from 2 to 4 on Monday and Wednesday with 113 and 114 teaching assistants and additional times as we find convenient. When the schedule is fixed, it will be circulated to all students in MA113 and 114. It is hoped that the extra work involved in holding office hours outside of the department will be balanced by having a quieter offices to work in the rest of the day. Pretest: Some (but not all) of our problems in calculus are due to the poor preparation of our students. Attached is a very simple pretest that you may give in the first recitation. This will help to warn the students that will be expected to know the basic skills from algebra in this course. The College of Engineering advisors are interested in identifying students who are not prepared and removing them from Calculus rather than failing them. Exams: There will be three exams and a final. These exams are scheduled in the evening at the times indicated in the common exam schedule. Though we have common exam times, we do not have common exams. Instructors in evening classes generally give their exams during a regularly scheduled class meeting. Rooms for exams will be assigned after classes begin. Instructors who wish to form coalitions to write common exams for several sections are encouraged to do so. If this is successful, I can arrange for the people involved to take over as coordinator for calculus. Homework: I expect that we will have undergraduate paper graders for MA113. The problems in the syllabus are a guide to the topics to be covered. Assignments to be graded should be taken from nearby even numbered problems. MA193: In addition, to the 4 hours of credit for MA113, the department offers one additional hour of credit as MA193 on a pass fail basis. Instructors may set their own grading policy for MA193. Typically, the grade is based on attendance in recitation with two unexcused absences allowed. Thus, a student with three or more unexcused absences will fail. Below are a few commons questions about MA193. Must a student take MA193? No, unless the student is in MathExcel. The math department recommends that all students take MA193. MathExcel students must take MA193. What section of MA193 should a student register in? Students should register in same section number for both MA193 and MA113. If a student drops or changes sections of MA113, they should also drop or change sections of MA193. Instructors and teaching assistants should check their MA193 rolls near the end of the semester. If you see a name that is unfamiliar, please try to determine if they are registered in another section of MA113. If you have a student who thinks they are registered in MA193, but does not appear on your roll, try to find out if they are registered in another section. Can MA193 be repeated? Yes, though there is little benefit to this. Can MA193 be taken without MA113? Yes, though it is not recommended. However, occasionally students will need to remain in MA193 in order to maintain a minimum number of credit hours for bureaucratic reasons. Suggested grading: Students need 90% for an A, 80% for a B, 70% for a C and 60% for a D. Grades should be based on three hour exams, quizzes, homework, and longer assignments.

3 hour exams  300
Final exam  100
Homework, quizzes, 150
TOTAL  550

Problems: Students should try to resolve problems with their instructor and/or teaching assistant. If this is not unsuccessful, the departmental ombudperson or the university ombudperson. If you have complaints about the syllabus, please let me know. Otherwise, you will be given the same syllabus next fall. Calculators: Most students are familiar with graphing calculators such as the TI-82. These calculators allow students to graph functions, solve equations, evaluate derivatives and definite integrals numerically. Elizabeth has TI-82 calculators that each instructor may check out. I suggest that students be allowed to use such calculators on exams. Test questions should be written so that it is clear whether a numerical answer from the calculator is acceptable or if students must carry out the computation by hand. In addition, some students will have machines that can carry out symbolic computations. I suggest that students not be allowed to use such machines on exams. Examples of such machines include the TI-89, TI-92, the HP48 and, of course, laptop computers. Computer labs: The math department has a computer lab, Inslab, that instructors may use. See http://www.ms.uky.edu/∼nslab for more information. This lab have Maple and Matlab available. The university also maintains numerous computer labs where students may use Maple and other mathematical software. These labs contain classrooms that may be reserved.

August 8, 2001